Given r = -0.90, MX = 4.13, SX = 1.77, MY = 3.45 and SY = 2.09, what is the regression equation?
Answer #1

Here it is given that 8 = -0.90 -0.90, Mx = 4.13, Sxal77 My = 3.45. Sy loog know the equation of regression line rsy (8) My S4.39+ 3.4 لا 27 ۰۵6 . = 00979 إليه و دو